OpenStudy (anonymous):

d to the 4th+32=12d squared? help please

jhonyy9 (jhonyy9):

so d to the 4th mean d with exponent 4 ?

OpenStudy (anonymous):

yes!

jhonyy9 (jhonyy9):

so than just note d^2 = t and will get an quadratic equation in t hope you can solve it

jhonyy9 (jhonyy9):

but dont forget it that t_1 and t_2 you need make equal d^2

jhonyy9 (jhonyy9):

t_1 and t_2 mean roots of quadratic equation in t ok ?

OpenStudy (jdoe0001):

\(\bf d^4+32=12d^2\implies d^4-12d^2+32=0\implies (d^{\color{red}{ 2}})^2-12(d^{\color{red}{2}})+32=0\) notice, tis just a quadratic

OpenStudy (jdoe0001):

shoot, got truncated... anyhow \(\bf d^4+32=12d^2\implies d^4-12d^2+32=0\\ \quad \\\implies (d^{\color{red}{ 2}})^2-12(d^{\color{red}{2}})+32=0\)